Find g'(x) if f(x)=|x-2| & g(x)=f[f(x)] for 2 less than or equal to x less than 4
A)1.   B)-1.   C)0.    D) none of these

Grade:12

2 Answers

seetaram dantu
7 years ago
for 2  x f(x) = x-2
 
now, g(x)= f[x-2] =f ([x] - 2)  we need to define piece wise  
1) for 2  x x
2) for 3  x
 
so g(x)=2 in case 1  and g(x)=1 in case 2  you can sustitute values of x for verification
 
 
seetaram dantu
7 years ago
keyboard errors:
case1 ==>, 2\leq x< 3, 2\leq x< 3case 1, 2\leq x< 3 =>g(x)= f(2-2)= f(0)= 2  as step x is 2 in this case  
 
case2==> 3\leq x< 4  g(x)= f(3-2)=f(1)=1  as step x is 3 in this case
 
sorry for the typing errors